# fastkqr: A Fast Algorithm for Kernel Quantile Regression

### v1

@@ E.4 Proof of Proposition 5
$\begin{split}\min_{\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}}\enskip{}&\sum^{T}_{t=1}\left[\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau}(y_{i}-b_{\tau_{t}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}})+\frac{\lambda_{2}}{2}\boldsymbol{\alpha}_{\tau_{t}}^{\top}\mathbf{K}\boldsymbol{\alpha}_{\tau_{t}}\right]\\ &+\lambda_{1}\sum^{T-1}_{t=1}V(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-b_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t+1}})\\ \text{subject to}\enskip{}&y_{i}=b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}},\,i\in S_{0,t},t=1,\ldots,T.\end{split}$ (50)
Therefore, $\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}$ is the minimizer of problem (50). Moreover, since $\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}$ is a feasible point of (50), we have
@@ E.5 Proof of Theorem 6
It then follows from (44) that
@@ E.6 Proof of Theorem 7
$\begin{split}&\min_{\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}}\enskip{}Q^{\gamma}\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}\right)\\ &\text{subject to}\enskip{}y_{i}=b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}},\,i\in\hat{S}_{t},1\leq t\leq T.\end{split}$ (51)
The Lagrangian of problem (51) is
$\begin{split}L(b,\boldsymbol{\alpha},\xi)=&\sum^{T}_{t=1}\left[\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,\tau}(y_{i}-b_{\tau_{t}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}})+\frac{\lambda_{2}}{2}\boldsymbol{\alpha}^{\top}_{\tau_{t}}\mathbf{K}\boldsymbol{\alpha}_{\tau_{t}}\right]\\ &+\sum^{T}_{t=1}\sum_{i\in\hat{S}_{t}^{\gamma}}\xi_{i,t}(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-y_{i})\\ &+\lambda_{1}\sum^{T-1}_{t=1}\sum_{i=1}^{n}V(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-b_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t+1}}),\end{split}$ (52)
where the $\xi_{i,t}$’s are the Lagrangian multipliers. Consider the following three cases, by the optimality of $\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}$ in problem (51),
@@
    $\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad+\lambda_{2}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t}=0,\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},\enskip{}i\in\hat{S}_{t}.\end{array}\right.$ (53)
@@
    $\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t}=0\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},\,i\in\hat{S}_{t}.\end{array}\right.$ (54)
@@
    $\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\sum_{i\in\hat{S}_{t}}\xi_{l,i}\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t}=0,\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},\,i\in\hat{S}_{t}.\end{array}\right.$ (55)
Therefore, it follows from expression (53) that
@@
$\left\{\begin{array}[]{l}\mathbf{0}=-\frac{1}{n}\sum_{i\notin\hat{S}_{t}}H^{\prime}_{\gamma,\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}-\frac{1}{n}\sum_{i\in\hat{S}_{t}}H^{\prime}_{\gamma,\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad\quad+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}\\ \quad\in-\frac{1}{n}\sum_{i\notin\hat{S}_{t}}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}-\frac{1}{n}\sum_{i\in\hat{S}_{t}}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad\quad+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}\\ \quad=-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i},\\ 0=-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\\ \quad\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t},\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},\,i\in\hat{S}_{t},\end{array}\right.$ (56)
Similarly, it can be shown for expressions (54) and (55), we have

### v2

@@ 7 Disclosure Statement
$\begin{split}\min_{\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}}\enskip{}&\sum^{T}_{t=1}\left[\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau}(y_{i}-b_{\tau_{t}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}})+\frac{\lambda_{2}}{2}\boldsymbol{\alpha}_{\tau_{t}}^{\top}\mathbf{K}\boldsymbol{\alpha}_{\tau_{t}}\right]\\ &+\lambda_{1}\sum^{T-1}_{t=1}V(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-b_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t+1}})\\ \text{subject to}\enskip{}&y_{i}=b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}},\,i\in S_{0,t},t=1,\ldots,T.\end{split}$ (S52)
Therefore, $\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}$ is the minimizer of problem (S52). Moreover, since $\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}$ is a feasible point of (S52), we have
@@
It then follows from (S46) that
@@
$\begin{split}&\min_{\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}}\enskip{}Q^{\gamma}\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}\right)\\ &\text{subject to}\enskip{}y_{i}=b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}},\,i\in\hat{S}_{t},1\leq t\leq T.\end{split}$ (S53)
The Lagrangian of problem (S53) is
$\begin{split}L(b,\boldsymbol{\alpha},\xi)=&\sum^{T}_{t=1}\left[\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,\tau}(y_{i}-b_{\tau_{t}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}})+\frac{\lambda_{2}}{2}\boldsymbol{\alpha}^{\top}_{\tau_{t}}\mathbf{K}\boldsymbol{\alpha}_{\tau_{t}}\right]\\ &+\sum^{T}_{t=1}\sum_{i\in\hat{S}_{t}^{\gamma}}\xi_{i,t}(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-y_{i})\\ &+\lambda_{1}\sum^{T-1}_{t=1}\sum_{i=1}^{n}V(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-b_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t+1}}),\end{split}$ (S54)
where the $\xi_{i,t}$’s are the Lagrangian multipliers. Consider the following three cases, by the optimality of $\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}$ in problem (S53),
@@
    $\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad+\lambda_{2}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t}=0,\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},\enskip{}i\in\hat{S}_{t}.\end{array}\right.$ (S55)
@@
    $\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t}=0\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},\,i\in\hat{S}_{t}.\end{array}\right.$ (S56)
@@
    $\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\sum_{i\in\hat{S}_{t}}\xi_{l,i}\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t}=0,\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},\,i\in\hat{S}_{t}.\end{array}\right.$ (S57)
Therefore, it follows from expression (S55) that
@@
$\left\{\begin{array}[]{l}\mathbf{0}=-\frac{1}{n}\sum_{i\notin\hat{S}_{t}}H^{\prime}_{\gamma,\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}-\frac{1}{n}\sum_{i\in\hat{S}_{t}}H^{\prime}_{\gamma,\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad\quad+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}\\ \quad\in-\frac{1}{n}\sum_{i\notin\hat{S}_{t}}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}-\frac{1}{n}\sum_{i\in\hat{S}_{t}}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad\quad+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}\\ \quad=-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i},\\ 0=-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\\ \quad\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t},\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},\,i\in\hat{S}_{t},\end{array}\right.$ (S58)
Similarly, it can be shown for expressions (S56) and (S57), we have